# A Survey on Dynamic Neural Networks: from Computer Vision to Multi-modal Sensor Fusion

### v3

@@ 2.1.2 Methods
Paper Main contribution Arch. Exit policy Application Section Görmez and Koyuncu (2024) Early Exits with threshold tailored to the class CNN & Transformer class confidence Image classification 2.1.3 Kim and Lee (2024) Method based on the Markov decision process for the early-exit decision CNN Learned policy Image classification 2.1.2 Liu et al. (2024c) Introduce an adapter for representations in a shared exit space Transf. Confidence Image classification 2.1.2 Meronen et al. (2024) Bayesian method to highlight out of distribution samples CNN - Image classification 2.1.2 Regol et al. (2024a) Combines quantization and Early Exit dynamic network Transf. Confidence Image classification 2.1.2 Regol et al. (2024b) Presents a loss for both accuracy and inference cost Transf. Confidence Image classification 2.1.2 Valade et al. (2024) The exiting is formalized as a classification with a reject option CNN Confidence Image classification 2.1.2 Yang et al. (2024b) Early Exits for object detection CNN Entropy Object detection 2.1.3 Zhang et al. (2024a) Self distillation with a dynamic generation of the importance weights CNN Confidence Image classification 2.1.2 Addad et al. (2025) Self-attention to improve feature representation and aggregation CNN - Image classification 2.1.1 Bajpai and Hanawal (2025) Reuse of previous exits responses for exit confidence CNN Confidence Image captioning 2.1.2 Kuhse et al. (2025) Early exit into YOLO architecture CNN - Object detection 2.1.3
@@ 2.1.3 Applications
Early Exits networks have been application in numerous computer vision tasks. In the context of object detection Yang et al. (2024b) proposes AdaDets, an architecture that speeds up inference through the use of Early Exits. The model formulates the bounding box prediction as a random variable following a certain probability distribution. This allows one to estimate the reliability of both the recognition and localization of a detected object. Furthermore, an entropy score is used for all the detections to estimate the certainty of the network toward the estimation, and make a decision on where to stop the inference. Recently, Kuhse et al. (2025) discusses a variation of the YOLO architecture to perform object detection with the option of exiting at early stages, in combination with specifically dedicated optimization algorithms.
@@ 2.2 Computational Routing
In Section 2.2.2, we illustrate various solutions for what concerns Mixture-of-Experts usage for image processing. Jacobs et al. (1991) introduces the original formulation of Mixture-of-Experts models. The key idea is to insert within a network a collection of expert subnetworks, which specialize in solving a subportion of the problem. This technique has been very popular recently with the advent of Large Language Models, where Mixture-of-Experts allows a network to scale the number of parameters while maintaining the computational cost constant. In this context, and differing from the original proposal, a gating module is introduced, which is in charge of selecting, among all available experts, a subportion of the most relevant ones to solve the task. The current use of MoE models differs from what has been presented in Jacobs et al. (1991): instead of having as result an ensemble of models, the most popular formulation use the experts as alternative layers activated according to the gating module. Formally, we can define a Mixture-of-Experts (MoE) model as a combination of the output of $N$ experts $f_{1},f_{2},...,f_{N}$ weighted by the gating function $g(x)$, producing the final output as $y=\sum_{i=1}^{N}g_{i}(x)f_{i}(x)$, where $g_{i}(x)$ represents the weight for the $i$-th expert and usually sums up to 1. Figure 1(b) provides a visual illustration of this technique.
@@
Similar to the Early Exits domain, numerical results for the publications in this section are not reported here. Also the field of Dynamic Routing is characterized by significant variations in test architectures and benchmark datasets, which prevents a meaningful direct comparison of performance. As previously discussed, this heterogeneity renders a fair quantitative analysis infeasible without extensive re-implementation and underscores the critical need for a standardized experimental protocol. A broader analysis of these limitations, and future directions to overcome these, are provided in the Discussion (Section 4).
@@ 2.2.2 Mixture of Experts
Odena et al. (2017) presents a very early proposal of a dynamic routing network composed of both MoE and Early Exits. The network allows a controller component to adaptively choose which expert to redirect to the next block, or it can decide to reduce computation and exit earlier the network. The controller is implemented as an Recurrent Neural Network by making the controller functions share weights and a hidden state, and it is trained using the REINFORCE algorithm (Williams, 1992). Gross et al. (2017) presents a CNN-based architecture in which all experts are first trained separately and then combined into a common decoder. Only one expert gets picked at each time according to a $k$-means clustering solution, which associates the choice of the expert classifier with a location in the feature space.
@@ 2.2.3 Applications
Pavlitskaya et al. (2020)
investigates a MoE architecture to achieve interpretability in the context of semantic segmentation while retaining comparable result quality. The principle consists in being able to observe both the overall model output as well as individual expert outputs, where the agreement or disagreement between those partial solutions can be used for insights into the decision process.
@@
Paper Summary Archit. Dynamic component Section Hazimeh et al. (2021) Address the problem of slow convergence in sparse gates CNN MoE 2.2.2 Li et al. (2021) Highlight the informative features in both the channel and spatial dimensions CNN Channel 2.2.1 Park et al. (2021) Adaptive routing path for Object Detection CNN Computational path 2.2.3 Riquelme et al. (2021) MoE in the context of ViT Transformer MoE 2.2.2 Sun et al. (2021) Adaptive model precision and frame skipping for Video Recognition CNN - 2.2.3 Wang et al. (2021a) Adaptive patch location for Video Recognition CNN Patch location 2.2.3 Liang et al. (2022a) Method to accelerate MoE for multi-task ViT Transformer MoE 2.2.2 Liu et al. (2022) Self-regulate computations according to performance target and resources available CNN Computational path 2.2.1 Meng et al. (2022) Block skipping in Transformer Transformer Block 2.2.1 Wang et al. (2022a) One stage adaptive patch location for Video Recognition CNN Patch location 2.2.3 Wang et al. (2022b) Adaptive patch location for Video Recognition with conditional exits CNN Patch location 2.2.3 Gao et al. (2023) Spatial and channel pruning hardware accelerator CNN Channel 2.2.3 Pan et al. (2023) Studies adversarial attacks on dynamic models CNN - 2.2.3 Verelst and Tuytelaars (2023) Adjusts dynamically the processing resolution of image regions CNN Computational path 2.2.1 Zhang et al. (2023a) Discuss adversarial attacks and robustness in the context of MoE CNN MoE 2.2.2 Devoto et al. (2025) Layer skipping in the fine-tuning of ViT Transformer Block 2.2.1 Liu et al. (2024a) Studies different approaches of MoE in the context of ViT Transformer MoE 2.2.2 Yang et al. (2024a) Adapts kernel weights and receptive field for different frames in Temporal Action Detection CNN Channel 2.2.3 Han et al. (2024) Empirical study on MoE design Transformer MoE 2.2.2 Jain et al. (2024) The model learns to dynamically process less important tokens with a less detailed representation Transformer MoE 2.2.2 Xie et al. (2024) Distillation among experts to allow each expert to get a better perception Transformer MoE 2.2.2 Huang et al. (2025) MoE applied to Point Cloud registration task Transformer MoE 2.2.3 Liu et al. (2025) Multiscale MoE for low latency and high accuracy in 3D object detection Transformer MoE 2.2.3 Sun et al. (2025) MoE strategy for reducing hyperspectral images computational complexity Transformer MoE 2.2.3
@@ 2.3 Token Skimming
In the context of sparsification of Transformer models for computation efficiency, skimming techniques consist of avoiding redundant computations by dynamically recognizing among all tokens the ones that are useless or redundant for the final task. In the case of ViTs, the assumption is that not all parts of an image contribute equally to solving the task, and many tokens contain irrelevant information. As an example, we can think of an image where most of the background is composed of a blue sky. Although the presence of a sky token can be semantically meaningful for a better understanding of the image context, the high number of tokens representing this information may harm the computational efficiency without necessary providing a better solution to the problem. The biological intuition is related to how humans efficiently perceive images and extract regions of interest from them (Zhou et al., 2017). By focusing only on the important information within a sequence, skimming helps to achieve faster inference speed and reduce noise within the model.
@@ 2.3.2 Token merging
Evo-ViT (Xu et al., 2022) is among the first to propose the merging of the unnecessary tokens together: the routing scheme divides the tokens in two parts where the top-$k$ important tokens are selected and go through the attention block (informative tokens), while the least important tokens are summarized and updated by a representative token (placeholder token), which then is also passed through the Transformer block. Centered Kernel Alignment (Kornblith et al., 2019) is used to measure the similarity of the intermediate token feature in each layer with respect to the classification token, starting from the assumption that the classification token can be used singularly since it is strongly correlated with classification. The informative tokens and the placeholder tokens are determined by a global class attention that gets updated across various layers.
@@
Wei et al. (2023) proposes a joint Token Pruning Squeezing (TPS) module. The idea is to have a scoring method based on Gumbel-Softmax to evaluate the tokens, which are then divided into two subsets. An alternative is also explored, where the class token attention values are used to measure the tokens’ importance—as already seen in EViT. The subset with pruned tokens is then fused into the subset with tokens reserved. An unidirectional nearest-neighbor matching algorithm is used to recognize the pruned tokens that are most similar to the reserved tokens. The pruned tokens are then fused with a weighting scheme based on the similarity of the reserved tokens. The similarity score used is the cosine similarity, while two solutions for placement of the TPS module are proposed: intrablocks (in between ViT blocks) and in-blocks (between MHSA and MLP). NEPAM (Jiang et al., 2023) performs token merging based on the similarity of pixel patches: tokens get divided into groups of neighbors, and the similarity of tokens within the group is then computed, using as a reference token the one in the upper-left corner of each group. The groups with the most similar tokens get merged. Furthermore, the method proposes a multi-scale relative position embedding based on the addition of a bias matrix to the tokens to solve the problem of corrupted position information of the merged embeddings.
@@ 3 Dynamic Sensor Fusion in the Context of Computer Vision
Analyzing the number of publications, there is clearly a research gap when compared to other applications of Adaptive Networks. On the other end, we can argue that Dynamic Neural Networks are very beneficial in the context of Sensor Fusion for several reasons. First, they serve efficiency purposes, and usually, Sensor Fusion applications run on low-computing devices or target real-time applications. Introducing some sort of elasticity in elaborating the different inputs allows for reduced computations. It enables to remove redundant and uninformative inputs and relates the use of certain sensors only to cases where the conditions benefit the overall prediction. In this sense, the benefit is derived from a more controlled (and tunable) trade-off between achieving a desired accuracy and constraining the computation resources. Furthermore, dynamically excluding certain sensors can help reduce the noise generated by the said sensors, in cases where the working conditions are suboptimal or a sensor failure is present. In these cases, a dynamic model can perform an improved selection of the optimal subset of sensors, taking into account also the environmental conditions (Chen et al., 2019a; Malawade et al., 2022b). Finally, these methods are able to identify and prioritize the elaboration of critical information by design. This is particularly useful in contexts where computation is a limited resource and the system needs to be responsive. We can imagine a setting where information can be elaborated either on-device or on the cloud. The introduced dynamicity in this case can help to better differentiate between information that is critical and needs to be computed on-premises for real-time and all-time responsiveness, and the information that can be sent out of the device to improve the overall model accuracy at the cost of latency or possibly information loss. In the following paragraphs, we present the application of Adaptive Sensor Fusion techniques for object detection, semantic segmentation, odometry, steering prediction and finally we introduce three different use cases of Sensor Fusion in the context of stress detection, power grid attack detection and multi-modal medical diagnosis.
@@
Mees et al. (2016) is the first to propose a novel adaptive fusion approach for object detection. The principle is to learn the weighting of predictions of different sensor modalities in an online manner with an architecture based on a CNN-based MoE. The model works with RGB, depth and motion information. The network performs late fusion: it uses different CNNs encoders for each modality and then a gating network which provides both a selection and the weighting of the experts outputs in order to get to the final classification values.
@@
Valada et al. (2017) propose AdapNet a dynamic architecture to perform semantic segmentation. It incorporates a convolutional MoE to dynamically fuse different modalities, where each expert is previously trained in a specific subset of the input space. Specifically, the paper proposes the fusion of RGB and depth modalities through the use of an adaptive gating network. Both modalities produce as intermediate output a segmentation mask. The gating mechanism then proceeds to act as a decision agent, which is allowed to pick one of the modalities output for each class in the segmentation task. The network is trained through a multi-stage training scheme, where the single experts are trained individually, and in a second phase, the gating parameters are also learned. Every dimensionality-reducing stage is split into convolutions applied at different scales based on ResNet.
@@
CARMA (Zhang et al., 2023b) is a framework proposing a fusion approach that uses the context to dynamically reconfigure the computation flow on a field-programmable gate array (FPGA) at runtime. This is performed through clock-gating unused sensors and model sub-components, with the aim of reducing the energy consumption. The architecture allows an optimization of the whole system, including sensors, model architecture, and hardware platform. A context identification process is enabled intermittently. The advantage of such a solution is to have global control over the overall power consumption of the whole system and not only the model part.

### v4

@@ 2.1.2 Methods
Paper Main contribution Arch. Exit policy Application Section Görmez and Koyuncu (2024) Early Exits with threshold tailored to the class CNN & Transformer class confidence Image classification 2.1.3 Kim and Lee (2024) Method based on the Markov decision process for the early-exit decision CNN Learned policy Image classification 2.1.2 Liu et al. (2024c) Introduce an adapter for representations in a shared exit space Transf. Confidence Image classification 2.1.2 Meronen et al. (2024) Bayesian method to highlight out of distribution samples CNN - Image classification 2.1.2 Regol et al. (2024a) Combines quantization and Early Exit dynamic network Transf. Confidence Image classification 2.1.2 Regol et al. (2024b) Presents a loss for both accuracy and inference cost Transf. Confidence Image classification 2.1.2 Valade et al. (2024) The exiting is formalized as a classification with a reject option CNN Confidence Image classification 2.1.2 Yang et al. (2024b) Early Exits for object detection CNN Entropy Object detection 2.1.3 Zhang et al. (2024a) Self distillation with a dynamic generation of the importance weights CNN Confidence Image classification 2.1.2 Addad et al. (2025) Self-attention to improve feature representation and aggregation CNN - Image classification 2.1.1 Bajpai and Hanawal (2025) Reuse of previous exits responses for exit confidence CNN Confidence Image captioning 2.1.2 Kuhse et al. (2025) Early exit into YOLO architecture CNN - Object detection 2.1.3 Wang et al. (2025) Adjust kernel shape based on the size of different objects CNN - Image enhancement 2.1.2
@@ 2.1.3 Applications
Early Exits networks have seen application in numerous computer vision tasks. In the context of object detection Yang et al. (2024b) proposes AdaDets, an architecture that speeds up inference through the use of Early Exits. The model formulates the bounding box prediction as a random variable following a certain probability distribution. This allows one to estimate the reliability of both the recognition and localization of a detected object. Furthermore, an entropy score is used for all the detections to estimate the certainty of the network toward the estimation, and make a decision on where to stop the inference. Recently, Kuhse et al. (2025) discusses a variation of the YOLO architecture to perform object detection with the option of exiting at early stages, in combination with specifically dedicated optimization algorithms.
@@ 2.2 Computational Routing
In Section 2.2.2, we illustrate various solutions regarding Mixture-of-Experts usage for image processing. Jacobs et al. (1991) introduces the original formulation of Mixture-of-Experts models. The key idea is to insert within a network a collection of expert subnetworks, which specialize in solving a subportion of the problem. This technique has been very popular recently with the advent of Large Language Models, where Mixture-of-Experts allows a network to scale the number of parameters while maintaining the computational cost constant. In this context, and differing from the original proposal, a gating module is introduced, which is in charge of selecting, among all available experts, a subportion of the most relevant ones to solve the task. The current use of MoE models differs from what has been presented in Jacobs et al. (1991): instead of resulting in an ensemble of models, the most popular formulation use the experts as alternative layers activated according to the gating module. Formally, we can define a Mixture-of-Experts (MoE) model as a combination of the output of $N$ experts $f_{1},f_{2},...,f_{N}$ weighted by the gating function $g(x)$, producing the final output as $y=\sum_{i=1}^{N}g_{i}(x)f_{i}(x)$, where $g_{i}(x)$ represents the weight for the $i$-th expert and usually sums up to 1. Figure 1(b) provides a visual illustration of this technique.
@@
Although the techniques detailed here are applied to visual processing components, Computational Routing shares deep roots and significant overlap with other deep learning domains. Most notably, the Mixture-of-Experts (MoE) paradigm has recently become unavoidable in the development of Large Language Models (LLMs). This mechanism allows a network to massively scale its parameters and capacity while maintaining a constant computational cost. This universal approach to solving efficiency bottlenecks means that advancements in routing policies or gating stability in vision seamlessly translate across the broader machine learning landscapes. Similar to the Early Exits domain, numerical results for the publications in this section are not reported here. Also the field of Dynamic Routing is characterized by significant variations in test architectures and benchmark datasets, which prevents a meaningful direct comparison of performance. As previously discussed, this heterogeneity renders a fair quantitative analysis infeasible without extensive re-implementation and underscores the critical need for a standardized experimental protocol. A broader analysis of these limitations, and future directions to overcome these, are provided in the Discussion (Section 4).
@@ 2.2.2 Mixture of Experts
Odena et al. (2017) presents a very early proposal of a dynamic routing network composed of both MoE and Early Exits. The network allows a controller component to adaptively choose which expert to redirect to the next block, or it can decide to reduce computation and exit earlier the network. The controller is implemented as an Recurrent Neural Network by making the controller functions share weights and a hidden state, and it is trained using the REINFORCE algorithm (Williams, 1992). Gross et al. (2017) presents a CNN-based architecture in which all experts are first trained separately and then combined into a common decoder. Only one expert gets picked at a time according to a $k$-means clustering solution, which associates the choice of the expert classifier with a location in the feature space.
@@ 2.2.3 Applications
Pavlitskaya et al. (2020) investigates a MoE architecture to achieve interpretability in the context of semantic segmentation while retaining comparable result quality. The principle consists in being able to observe both the overall model output as well as individual expert outputs, where the agreement or disagreement between those partial solutions can be used for insights into the decision process.
@@
Paper Summary Archit. Dynamic component Section Hazimeh et al. (2021) Address the problem of slow convergence in sparse gates CNN MoE 2.2.2 Li et al. (2021) Highlight the informative features in both the channel and spatial dimensions CNN Channel 2.2.1 Park et al. (2021) Adaptive routing path for Object Detection CNN Computational path 2.2.3 Riquelme et al. (2021) MoE in the context of ViT Transformer MoE 2.2.2 Sun et al. (2021) Adaptive model precision and frame skipping for Video Recognition CNN - 2.2.3 Wang et al. (2021a) Adaptive patch location for Video Recognition CNN Patch location 2.2.3 Liang et al. (2022a) Method to accelerate MoE for multi-task ViT Transformer MoE 2.2.2 Liu et al. (2022) Self-regulate computations according to performance target and resources available CNN Computational path 2.2.1 Meng et al. (2022) Block skipping in Transformer Transformer Block 2.2.1 Wang et al. (2022a) One stage adaptive patch location for Video Recognition CNN Patch location 2.2.3 Wang et al. (2022b) Adaptive patch location for Video Recognition with conditional exits CNN Patch location 2.2.3 Gao et al. (2023) Spatial and channel pruning hardware accelerator CNN Channel 2.2.3 Pan et al. (2023) Studies adversarial attacks on dynamic models CNN - 2.2.3 Verelst and Tuytelaars (2023) Adjusts dynamically the processing resolution of image regions CNN Computational path 2.2.1 Zhang et al. (2023a) Discuss adversarial attacks and robustness in the context of MoE CNN MoE 2.2.2 Devoto et al. (2025) Layer skipping in the fine-tuning of ViT Transformer Block 2.2.1 Liu et al. (2024a) Studies different approaches of MoE in the context of ViT Transformer MoE 2.2.2 Yang et al. (2024a) Adapts kernel weights and receptive field for different frames in Temporal Action Detection CNN Channel 2.2.3 Han et al. (2024) Empirical study on MoE design Transformer MoE 2.2.2 Jain et al. (2024) The model learns to dynamically process less important tokens with a less detailed representation Transformer MoE 2.2.2 Xie et al. (2024) Distillation among experts to allow each expert to get a better perception Transformer MoE 2.2.2 Huang et al. (2025) MoE applied to Point Cloud registration task Transformer MoE 2.2.3 Liu et al. (2025) Multiscale MoE for low latency and high accuracy in 3D object detection Transformer MoE 2.2.3 Sun et al. (2025) MoE strategy for reducing hyperspectral images computational complexity Transformer MoE 2.2.3 Chen et al. (2025) Chen et al. (2025) MoE framework for Vision Language medical diagnosis Transformer MoE 2.2.3
@@ 2.3 Token Skimming
In the context of sparsification of Transformer models for computation efficiency, skimming techniques consist of avoiding redundant computations by dynamically recognizing among all tokens the ones that are useless or redundant to the final task. In the case of ViTs, the assumption is that not all parts of an image contribute equally to solving the task, and many tokens contain irrelevant information. As an example, we can think of an image where most of the background is composed of a blue sky. Although the presence of a sky token can be semantically meaningful for a better understanding of the image context, the high number of tokens representing this information may harm the computational efficiency without necessary providing a better solution to the problem. The biological intuition is related to how humans efficiently perceive images and extract regions of interest from them (Zhou et al., 2017). By focusing only on the important information within a sequence, skimming helps to achieve faster inference speed and reduce noise within the model.
@@
Among all presented methods, Token Skimming is the most reusable across domains. This is primarily due to the underlying Transformer architecture treating inputs as sets of separate tokens that interact with each other in the self-attention block, regardless of what those tokens represent. This allows the algorithms for Token Skimming to be agnostic to the data’s origin. Exceptions exist where the skimming process relies explicitly on 2D spatial properties; for instance, methods that utilize spatial adjacency to determine propagation graphs or group neighboring pixel patches. These methods restrict their direct application to non-visual tasks without some structural adaptations.
@@ 2.3.2 Token merging
Evo-ViT (Xu et al., 2022) is among the first to propose the merging of the unnecessary tokens together: the routing scheme divides the tokens into two parts where the top-$k$ important tokens are selected and go through the attention block (informative tokens), while the least important tokens are summarized and updated by a representative token (placeholder token), which then is also passed through the Transformer block. Centered Kernel Alignment (Kornblith et al., 2019) is used to measure the similarity of the intermediate token feature in each layer with respect to the classification token, starting from the assumption that the classification token can be used individually since it is strongly correlated with classification. The informative tokens and the placeholder tokens are determined by a global class attention that gets updated across various layers.
@@
Wei et al. (2023) proposes a joint Token Pruning Squeezing (TPS) module. The idea is to have a scoring method based on Gumbel-Softmax to evaluate the tokens, which are then divided into two subsets. An alternative is also explored, where the class token attention values are used to measure the tokens’ importance—as already seen in EViT. The subset with pruned tokens is then fused into the subset with tokens reserved. An unidirectional nearest-neighbor matching algorithm is used to identify the pruned tokens that are most similar to the reserved tokens. The pruned tokens are then fused with a weighting scheme based on the similarity of the reserved tokens. The similarity score used is the cosine similarity, while two solutions for placement of the TPS module are proposed: intrablocks (in between ViT blocks) and in-blocks (between MHSA and MLP). NEPAM (Jiang et al., 2023) performs token merging based on the similarity of pixel patches: tokens get divided into groups of neighbors, and the similarity of tokens within the group is then computed, using as a reference token the one in the upper-left corner of each group. The groups with the most similar tokens get merged. Furthermore, the method proposes a multi-scale relative position embedding based on the addition of a bias matrix to the tokens to solve the problem of corrupted position information of the merged embeddings.
@@ 3 Dynamic Sensor Fusion in the Context of Computer Vision
Analyzing the number of publications, there is clearly a research gap when compared to other applications of Adaptive Networks. On the other hand, we can argue that Dynamic Neural Networks are very beneficial in the context of Sensor Fusion for several reasons. First, they serve efficiency purposes, and usually, Sensor Fusion applications run on low-computing devices or target real-time applications. Introducing some sort of elasticity in elaborating the different inputs allows for reduced computations. It enables to remove redundant and uninformative inputs and relates the use of certain sensors only to cases where the conditions benefit the overall prediction. In this sense, the benefit is derived from a more controlled (and tunable) trade-off between achieving a desired accuracy and constraining the computation resources. Furthermore, dynamically excluding certain sensors can help reduce the noise generated by the said sensors, in cases where the working conditions are suboptimal or a sensor failure is present. In these cases, a dynamic model can perform an improved selection of the optimal subset of sensors, taking into account also the environmental conditions (Chen et al., 2019a; Malawade et al., 2022b). Finally, these methods are able to identify and prioritize the elaboration of critical information by design. This is particularly useful in contexts where computation is a limited resource and the system needs to be responsive. We can imagine a setting where information can be elaborated either on-device or on the cloud. The introduced dynamicity in this case can help to better differentiate between information that is critical and needs to be computed on-premises for real-time and all-time responsiveness, and the information that can be sent out of the device to improve the overall model accuracy at the cost of latency or possibly information loss.
When it comes to Dynamic Neural Networks for Sensor Fusion the field remains relatively underexplored compared to unimodal Computer Vision. Furthermore, the deployment of dynamic multi-modal architectures introduces specific failure modes, operational constraints, and evaluation bottlenecks that require further investigation. The primary vulnerability of dynamic fusion architectures lies in the gating or routing mechanisms. Gating modules are usually trained to assess sensor reliability based on the environmental context. In cases of Out-of-Distribution (OOD) conditions or extreme domain shifts, the gating network can become miscalibrated. This leads to a failure in identifying corrupted modalities, routing noisy data into the computation path. Furthermore, the dynamic router can disproportionately favor a high-bandwidth sensor (e.g., RGB camera) during the training phase. This leads to modality dominance, effectively reducing the dynamic fusion to a static unimodal inference path and neutralizing the benefits of the multi-sensor setup. Lastly, Sensor Fusion involves dealing with sensor synchronization. The architecture has to process asynchronous data streams operating at heterogeneous frequencies (e.g., cameras at 30Hz, LiDAR at 10Hz, IMU at 200Hz). While dynamic networks theoretically offer the ability to adaptively handle variable-rate streams, synchronizing these inputs dynamically on the fly introduces significant computational overhead. In latency-critical applications such as autonomous navigation, the temporal cost of evaluating the gating function to select the optimal sensor subset is not completely clear and could potentially negate the latency saved by skipping computations. Furthermore, in distributed systems where computation is offloaded to edge units, fluctuating communication bandwidth and transmission latency need to be taken into account on top of theoretical FLOP reduction.
Finally, it is clear that the field is missing proper evaluation tests to correctly identify the robustness in case of missing or degraded modalities. It is important to highlight the behavior of the model under varying degrees of sensor noise, inducing random modality dropouts, and testing under artificial latency constraints. These types of measurements are fundamental for estimating the network’s ability to operate under safety requirements in case of individual or multiple sensor failures. Generally speaking, the transition toward practical applicability and rigorous real-world testing remains a missing step in the current research.
In the following paragraphs, we present the application of Adaptive Sensor Fusion techniques for object detection, semantic segmentation, odometry, steering prediction and finally we introduce three different use cases of Sensor Fusion in the context of stress detection, power grid attack detection and multi-modal medical diagnosis.
@@
Mees et al. (2016) is the first to propose a novel adaptive fusion approach for object detection. The principle is to learn the weighting of predictions of different sensor modalities in an online manner with an architecture based on a CNN-based MoE. The model works with RGB, depth and motion information. The network performs late fusion: it uses different CNN encoders for each modality and then a gating network which provides both a selection and the weighting of the experts outputs in order to get to the final classification values.
@@
Valada et al. (2017) propose AdapNet a dynamic architecture to perform semantic segmentation. It incorporates a convolutional MoE to dynamically fuse different modalities, where each expert is previously trained in a specific subset of the input space. Specifically, the paper proposes the fusion of RGB and depth modalities through the use of an adaptive gating network. Both modalities produce as intermediate output a segmentation mask. The gating mechanism then proceeds to act as a decision agent, which is allowed to pick one of the modality outputs for each class in the segmentation task. The network is trained through a multi-stage training scheme, where the single experts are trained individually, and in a second phase, the gating parameters are also learned. Every dimensionality-reducing stage is split into convolutions applied at different scales based on ResNet.
@@
CARMA (Zhang et al., 2023b) is a framework proposing a fusion approach that uses the context to dynamically reconfigure the computation flow on a field-programmable gate array (FPGA) at runtime. This is performed by clock-gating unused sensors and model sub-components, with the aim of reducing the energy consumption. The architecture allows an optimization of the whole system, including sensors, model architecture, and hardware platform. A context identification process is enabled intermittently. The advantage of such a solution is to have global control over the overall power consumption of the whole system and not only the model part.